OpenStudy (anonymous):

Suppose 0 < A < AB < 1 < B Rank the following five expressions from least to greatest by selecting the correct order of the expressions in the inequality. 1) 0 2) Log(A) 2) Log(b) 3) Log(A)+Log(B) 5) Log(B^A)

ganeshie8 (ganeshie8):

I dont see any difference between 3rd and 4th expressions ?

OpenStudy (anonymous):

Sorry, my bad! Does that make more sense?

ganeshie8 (ganeshie8):

Since A and AB are less than 1, log(A) and log(AB) will be negative

ganeshie8 (ganeshie8):

see if u can sort them ..

OpenStudy (anonymous):

What are the rules for determining positive/negative based on <=>1?

OpenStudy (anonymous):

Like I get that a and a*b are gunna be fractions but I dont understand once their put into a log how you can tell what they will come out to be?

ganeshie8 (ganeshie8):

say \(a \lt 1\), consider below : \(\log_{10} (a)\)

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

look at the graph, whats the value of `log` function when x is less than 1 ?

ganeshie8 (ganeshie8):

if u dont like graphs : \(\log_{10} (a) = x \implies a = 10^x\) since \(a\) is a fraction, \(x\) needs to be a negative (why ?)

OpenStudy (tkhunny):

\(1 < B^{A} < B\) This makes: \(0 < A < AB < 1 < B^{A}< B \) The logarithm function preserves order. This is a property that makes it such a valuable transformation in many fields and exercises. Thus: \(log(A) < log(AB) < log(1) < log(B^{A})< log(B)\) -- We have to lose the zero (0). A little logarithm magic, gives: \(log(A) < log(AB) = log(A) + log(B) < log(1) = 0 < log(B^{A}) = A\cdot log(B)< log(B)\) Note: It is generally considered bad form to change variables in the middle of a problem. Don't start talking about "B" and then end up talking about "b".

OpenStudy (anonymous):

......

ganeshie8 (ganeshie8):

-_-_-

ganeshie8 (ganeshie8):

ask a question if u dont understand lol.. hard to understand dot language :P

OpenStudy (anonymous):

That's because I have no idea what to do with the information i was just spewed at lol...I think we should just redirect back to your version lol

OpenStudy (tkhunny):

There is not much there. 1) Put \(B^{A}\) in the list. 2) Apply the logarithm function. 3) Use a few logarithm rules. Look at each piece. Let it soak in.

ganeshie8 (ganeshie8):

first things first : you need to know when a log value is negative, and when it is positive

ganeshie8 (ganeshie8):

did u get when it would be negative/positive ?

OpenStudy (anonymous):

I understand in exponential form but I guess I'm just having difficulty applying that to log form

ganeshie8 (ganeshie8):

okay, the graph was not useful is it ?

OpenStudy (anonymous):

no not really, we didn't spend much time on graphing

ganeshie8 (ganeshie8):

okay, then discard the graph lets see how to figure out the log value from exponent perspective

ganeshie8 (ganeshie8):

\(\large \log_{10} (a) = x ~~~~ \implies~~~~ a = 10^x\)

ganeshie8 (ganeshie8):

fine with this right ?

OpenStudy (anonymous):

yeah, and since a is a fraction that would make x negative eg Log(a)=-# correct?

ganeshie8 (ganeshie8):

you got it !

ganeshie8 (ganeshie8):

\(a = 10^x\) for example, when \(x = -1\) : \(a = 10^{-1} = 0.1\) which is a fraction !

ganeshie8 (ganeshie8):

so exponent has to be negative, for the value to evaluate to a fraction

ganeshie8 (ganeshie8):

log value = exponent = x